It would take an qpprentice house painter 1.5 h longer than his supervisor to paint an apartment. If they work together, they can complete the job in 4 h.
About how long would it take the apprentice to complete the job working alone? Round your answer to the nearest tenth of an hour.

Answer by richwmiller(17219)

1/x+1/(x+1.5)=1/4
x = 7.31971
x=7.3 for the supervisor
x+1.5=8.8 hours for the apprentice.